OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4,1) with slope -4.

OpenStudy (anonymous):

(y -y1) = m(x-x1) + c m = -4 x1,y1 = 4,1

OpenStudy (igreen):

Point Slope Form = \(y - y_1 = m(x - x_1)\) Where \(y_1\) = y-value Where \(x_1\) = x-value Where \(m\) = slope

OpenStudy (igreen):

Can you plug the terms in now? @Kahlea_renee

OpenStudy (anonymous):

Thank you yes I can

OpenStudy (igreen):

No problem.